bemine11:

Solve -12 = z ÷ 3 for z. z =

bemine11:

would it be -36

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bemine11 would it be -36 \(\color{#0cbb34}{\text{End of Quote}}\) Yes

bemine11:

thanks

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bemine11 thanks \(\color{#0cbb34}{\text{End of Quote}}\) np

bemine11:

have a good day guys

xxDeppressionxx:

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : z/12-(3/z)=0 3 Simplify — z z 3 —— - — = 0 12 z z Simplify —— 12 z 3 —— - — = 0 12 z Find the Least Common Multiple The left denominator is : 12 The right denominator is : z Number of times each prime factor appears in the factorization of: Prime Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} 2 2 0 2 3 1 0 1 Product of all Prime Factors 12 1 12 Number of times each Algebraic Factor appears in the factorization of: Algebraic Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} z 0 1 1 Least Common Multiple: 12z Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator z • z - (3 • 12) z2 - 36 ———————————————— = ——————— 12z 12z Factoring: z2 - 36 Factoring: z2 - 36 Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2 - AB + BA - B2 = A2 - AB + AB - B2 = A2 - B2 Note : AB = BA is the commutative property of multiplication. Note : - AB + AB equals zero and is therefore eliminated from the expression. Check : 36 is the square of 6 Check : z2 is the square of z1 Factorization is : (z + 6) • (z - 6) (z + 6) • (z - 6) ——————— = 0 12z When a fraction equals zero ... Where a fraction equals zero, its numerator, the part which is above the fraction line, must equal zero. Now,to get rid of the denominator, Tiger multiplys both sides of the equation by the denominator. Here's how: (z+6)•(z-6) ——————————— • 12z = 0 • 12z 12z Now, on the left hand side, the 12z cancels out the denominator, while, on the right hand side, zero times anything is still zero. The equation now takes the shape : (z+6) • (z-6) = 0 A product of several terms equals zero. When a product of two or more terms equals zero, then at least one of the terms must be zero. We shall now solve each term = 0 separately In other words, we are going to solve as many equations as there are terms in the product Any solution of term = 0 solves product = 0 as well. Solve : z+6 = 0 Subtract 6 from both sides of the equation : z = -6 Solve : z-6 = 0 Add 6 to both sides of the equation : z = 6 z = 6 z = -6

xxDeppressionxx:

Supie it is not 36 because it wants to know what z is

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx Supie it is not 36 because it wants to know what z is \(\color{#0cbb34}{\text{End of Quote}}\) Yes it wants to know what z is and ik that it's not 36

xxDeppressionxx:

And Z is 6

supie:

No it's not.

xxDeppressionxx:

What is it then can you explain what it is?

supie:

The answer is -36.

xxDeppressionxx:

Yes

supie:

You said that it was 6 \(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx And Z is 6 \(\color{#0cbb34}{\text{End of Quote}}\)

xxDeppressionxx:

I know i said it was 6 but i forgot to add the 3 and the negative

supie:

Ok then I was right-

xxDeppressionxx:

Yes